OpenStudy (anonymous):

what is 5% of 80? plz show work

OpenStudy (anonymous):

5% means 5/100 hence 5% of 80=5/100 X 80= 4

OpenStudy (anonymous):

5x80/100

OpenStudy (anonymous):

\[5 \% \; \; of \; \; 80\] \[\implies \frac{5}{100} \times 80 = ??\] Solve this and tell me what are you getting??